# Emilia: A Large-Scale, Extensive, Multilingual, and Diverse Dataset for Speech Generation

### v1

@@ Abstract
Recent advancements in speech generation have been driven by the large-scale training datasets. However, current models fall short of capturing the spontaneity and variability inherent in real-world human speech, due to their reliance on audiobook datasets limited to formal read-aloud speech styles. To bridge this gap, we introduce Emilia-Pipe, an open-source preprocessing pipeline to extract high-quality training data from valuable yet underexplored in-the-wild data that capture spontaneous human speech in real-world contexts. By leveraging Emilia-Pipe, we construct Emilia, the first multilingual speech generation dataset derived from in-the-wild speech data. This dataset comprises over 101k hours of speech across six languages: English, Chinese, German, French, Japanese, and Korean. Besides, we expand Emilia to Emilia-Large, a dataset exceeding 216k hours, making it the largest open-source speech generation dataset available. Extensive experiments demonstrate that Emilia significantly outperforms traditional audiobook datasets in generating spontaneous and human-like speech, showcasing superior performance in capturing diverse speaker timbre and speaking styles of real-world human speech. Furthermore, this work underscores the importance of scaling dataset size to advance speech generation research and validates the effectiveness of Emilia for both multilingual and crosslingual speech generation.
@@
Speech Generation, TTS, Dataset, Multilingual, In-the-wild data
@@ I Introduction
In recent years, (zero-shot) speech generation research has witnessed significant advancements [1, 2, 3, 4, 5, 6, 7], with diverse models utilizing large-scale training datasets. These advancements lead to improved voice quality, timbre similarity, and naturalness [8]. Nevertheless, generated speech still falls short of replicating the spontaneity and variability characteristic of real-world human speech [2, 9].
@@
A primary factor for this limitation is the reliance of current speech generation models on datasets derived from audiobooks [10, 11]. Such datasets predominantly feature formal, read-aloud speech styles, which starkly contrast with the diverse and spontaneous nature of human speech in casual or conversational settings. Real-world speech is characterized by a wide range of phenomena, including breathing sounds, pauses, stuttering, repetitions, variations in speaking rate and emotions. Therefore, there is a significant research gap for a new dataset that captures a broader spectrum of speech styles to advance the field toward generating more spontaneous and human-like speech.
@@
However, directly utilizing in-the-wild speech data presents significant challenges due to variations in quality, typically manifested as frequent background noise or music, reverberation, overlapping speakers within a single sample, inconsistent audio lengths, and the absence of essential annotations such as text transcriptions [12, 13]. Training speech generation models on such unprocessed raw data can lead to degraded performance [12, 14, 15, 16]. Although prior studies [12, 13] have proposed automatic preprocessing pipelines to address challenges associated with in-the-wild speech data, these pipelines heavily rely on proprietary models, significantly limiting their accessibility to the broader research community. Additionally, these pipelines are restricted to monolingual (i.e., Chinese-only) speech data, rendering them unsuitable for processing the vast multilingual speech data available in the wild. Furthermore, the computational efficiency of these pipelines remains undocumented, raising concerns about their practicality for scaling speech generation datasets. There is an urgent need for an effective open-source preprocessing pipeline capable of handling multilingual in-the-wild speech data efficiently and facilitating substantial dataset expansion.
@@
In response, we introduce Emilia-Pipe, the first open-source preprocessing pipeline designed to harness valuable yet underexplored in-the-wild multilingual speech data for building high-quality training datasets of spontaneous and human-like speech generation models. The Emilia-Pipe is designed to comprise six core preprocessing steps: standardization, source separation, speaker diarization, fine-grained segmentation via voice activity detection (VAD), automated speech recognition (ASR), and filtering. Furthermore, Emilia-Pipe incorporates extensive engineering optimizations to enhance both robustness and efficiency. These designs position Emilia-Pipe as a versatile and scalable tool for constructing extensive, multilingual, and diverse speech datasets, thereby addressing the critical gap in speech generation. Leveraging Emilia-Pipe, we construct Emilia, the first multilingual speech generation dataset derived from in-the-wild speech data, which comprises over 101k hours of speech data at 24 kHz and spans six languages: English (En), Chinese (Zh), German (De), French (Fr), Japanese (Ja), and Korean (Ko). Additionally, we further expand the Emilia dataset to an even larger scale of 216K hours, introducing Emilia-Large. Table I compares Emilia and Emilia-Large with several existing speech generation datasets. The key advantages of the Emilia and Emilia-Large datasets are summarized as follows.
@@
  • Extensive: The Emilia dataset contains over 101k hours of speech data at 24 kHz, and the Emilia-Large variant expands the size to 216k hours. To our best, Emilia-Large is the largest open-source speech generation dataset.
@@
  • Multilingual: The Emilia and Emilia-Large datasets covers six languages, supporting the training of multilingual and crosslingual speech generation models.
@@
  • Diverse: The most significant feature of Emilia and Emilia-Large, compared to previous datasets, is its composition of mostly spontaneous, in-the-wild speech data, encompassing a wide range of speaking styles, which is crucial for training the next generation of speech generation models capable of producing spontaneous and human-like speech.
@@
  • Dynamic: The Emilia and Emilia-Large dataset uniquely features an automatic and efficient processing pipeline, i.e., Emilia-Pipe, enabling seamless expansion in total size and language coverage, significantly facilitating rapid and scalable dataset creation.
@@
Libri-Heavy [22] Audiobook 50k En 16k
MLS [11] Audiobook 51k En/Fr/De/Nl/Es/It/Pt/Pl 16k
Libri-Light [10] Audiobook 60k En 16k
@@
  • Larger-Scale Dataset: We expanded the initial Emilia dataset to create Emilia-Large, a dataset more than twice the size of its predecessor, standing as the largest open-source speech generation dataset available to date.
  • Comparative Analysis of Audiobook and In-the-Wild Datasets: We present more details for the experiments that compare the speech generation performance of audiobook datasets and the in-the-wild dataset Emilia.
  • Exploration of Data Scaling Laws in Speech Generation: Leveraging the large-scale nature of Emilia, we conducted experiments to investigate the impact of dataset size on speech generation performance. These experiments provide valuable insights into the importance of scaling dataset size for advancing the field of speech generation.
  • Multilingual and Crosslingual Effectiveness Analysis: We performed a more comprehensive set of experiments to validate the multilingual and crosslingual capabilities of the Emilia dataset, further demonstrating its versatility and applicability across diverse linguistic contexts.
Our code for the Emilia-Pipe [^1] https://github.com/open-mmlab/Amphion/tree/main/preprocessors/Emilia and the Emilia and Emilia-Large datasets [^2] https://huggingface.co/datasets/amphion/Emilia-Dataset has been made publicly available to facilitate further research and reproducibility. Additionally, during the course of this work, significant advancements in speech generation research have emerged using the initial Emilia dataset. These include, but are not limited to, MaskGCT [24], F5-TTS [25], and Vevo [26], all of which provide open-source pre-trained models or code, serving as valuable contributions to the research community. The remainder of this paper is structured as follows. Section II reviews related work, followed by a comprehensive description of the proposed Emilia-Pipe in Section III. Section IV details and analyzes the construction of the Emilia dataset. Implementation specifics and experimental results are presented in Section V, accompanied by an in-depth discussion. Finally, Section VI concludes the paper.
@@ II-A Speech Generation Datasets
The size of datasets for speech generation has increased substantially over the years. Early datasets typically comprised tens of hours of speech. For example, LJSpeech [17] contains 24 hours of speech data from a single speaker. The VCTK dataset [18] includes 44 hours of speech data from 109 speakers, while AISHELL-3 [19] comprises approximately 85 hours of recordings from 218 speakers. Subsequently, larger-scale datasets have emerged to enable research in zero-shot speech generation. For instance, LibriTTS [20] contains 585.8 hours of speech data from audiobooks, while GigaSpeech offers 10k hours of unprocessed speech data from diverse sources in the wild. Later, the research community scaled up speech generation datasets to over 50k hours. Emerging datasets such as MLS [11] (51k hours), Libri-Light [10] (60k hours), and Libri-Heavy [22] (50k hours) significantly enhance the performance for zero-shot speech generation models. However, these datasets are either rooted in audiobooks, limited to formal read-aloud speech styles due to their reliance on audiobooks as data sources, which lack the naturalness and variability of spontaneous speech [9, 27], or they directly utilize raw in-the-wild speech data that suffers from uncertain quality [12, 14, 15, 16]. Both scenarios lead to sub-optimal performance in speech generation tasks.
Notably, two previous works [13, 12] propose similar automated preprocessing pipelines for building speech generation datasets from in-the-wild data. However, as discussed earlier, these pipelines heavily rely on proprietary models and are restricted to Chinese-only speech data, with unknown computational efficiency. To bridge this gap, we design and open-source an efficient pipeline, Emilia-Pipe, which can rapidly process large-scale raw multilingual speech data to facilitate substantial dataset creation and expansion.
@@ II-B Speech Generation Models
Traditional speech generation models, such as Tacotron [28], FastSpeech [29, 30], and NaturalSpeech [27], are limited by smaller datasets comprising tens of hours of speech data and a limited number of speakers. Recent advancements in speech generation, driven by large-scale speech datasets, have significantly improved voice quality, timbre similarity, and naturalness in zero-shot speech generation using only a short reference speech sample of a few seconds [1, 2, 3, 4, 5, 6]. For example, VALLE [1], VoiceBox [3], and SoundStorm [4] use more than 50k hours of speech data for training. Later models, such as NaturalSpeech3 [2], BaseTTS [5], and Seed-TTS [6], further expanded dataset sizes to over 100k hours. Notably, BaseTTS reported the “emergent abilities” of TTS models as dataset sizes scaled, enabling the rendering of complex prosody patterns such as emotions based on textual cues without explicit labels. An large-scale open-source dataset is needed for large-scale speech generation research. This work addresses this specific need.
Figure 1: An overview of the Emilia-Pipe processing pipeline. It consists of six steps, namely, standardization, source separation, speaker diarization, fine-grained segmentation by voice activity detection (VAD), automated speech recognition (ASR), and filtering.
@@ III The Emilia-Pipe Processing Pipeline
As discussed above, raw speech data in the wild need to be processed to be leveraged for training speech generation models. Therefore, we design an automatic preprocessing pipeline, Emilia-Pipe, for transforming in-the-wild multilingual speech data into high-quality training datasets. As illustrated in Fig. 1, Emilia-Pipe includes six steps: standardization, source separation, speaker diarization, fine-grained segmentation by voice activity detection (VAD), automated speech recognition (ASR), and filtering. This section details the six steps of Emilia-Pipe and evaluates its performance.
@@
The raw speech data in the wild vary in encoding formats, sampling rates, and other characteristics. To standardize the collected data, we convert all samples to WAV files, set them to a mono channel, and resample to 24 kHz. We set the sample width to 16 bits and adjust the target decibels relative to full scale to -20 dBFS. The actual gain is constrained within -3 to 3 dB to ensure appropriate volume without distortion. Finally, we normalize the waveform by dividing each sample by the maximum amplitude, ensuring values range between -1 and 1. This step ensures a consistent data format for further processing.
@@
The raw speech data in the wild often contain background music and noise, which negatively impact the performance of speech generation models [3, 14, 15]. To address this issue, we employ source separation techniques to extract clean human vocals. Specifically, we utilize the open-source pre-trained Ultimate Vocal Remover model [31], UVR-MDX-Net Inst 3333https://github.com/TRvlvr/model_repo/releases/tag/all_public_uvr_models. This model achieves a high signal-to-distortion ratio (SDR) of 11.15 for vocal separation on the Synth MVSep dataset [32]. Using this model, we effectively separate human vocals for further processing.
@@
After extracting clean human vocals from the raw speech data, we apply speaker diarization techniques to partition long-form speech data into multiple utterances based on the speaker. This process generates a series of speech segments, with each segment ideally containing only one speaker, ensuring compatibility with existing datasets for speech generation. To achieve this, we leverage the PyAnnote speaker diarization 3.1 pipeline. [^4] https://github.com/pyannote/pyannote-audio, which includes three core components: speaker segmentation, speaker embedding, and clustering, achieving state-of-the-art (SOTA) speaker diarization performance [33, 34]. The output is a list of temporal annotations indicating the start and end times of the single-speaker segments.
@@
Although the speaker diarization pipeline provides a coarse segmentation for the raw speech data, the resulting utterances may still be too long to fit into memory. To address this, we use a VAD model to further segment the utterances into smaller segments ranging from 3 to 30 seconds. This is achieved by concatenating consecutive chunks containing voice activity from the same speaker. We leverage the open-source library Silero-VAD [^5] https://github.com/snakers4/silero-vad. The pre-trained model provided in Silero-VAD achieves a ROC-AUC score of 0.99 on the LibriParty dataset, ensuring accurate detection of voice activity. [^6] https://github.com/speechbrain/speechbrain/tree/develop/recipes/LibriParty/generate_dataset
@@
The absence of text transcriptions impedes the direct use of in-the-wild dataset for TTS. Therefore, we use ASR techniques to transcribe the segmented speech data. Considering the trade-off among speed, robustness, and accuracy, we employ the medium version of the Whisper model [35], a SOTA multilingual ASR model capable of robust speech translation and language identification. To further enhance efficiency, we leverage WhisperX [36], which builds on the faster-whisper backend [^7] https://github.com/SYSTRAN/faster-whisper and the CTranslate2 engine. [^8] https://github.com/OpenNMT/CTranslate2 This setup is up to four times faster than the official Whisper implementation while maintaining the comparable accuracy and using less memory. Additionally, we omit the original model’s inherent VAD component by using the outputs in the last step to avoid redundant processing and develop batched inference to transcribe the speech data in parallel. These improvements allow our ASR step to achieve accurate text transcriptions for the speech data with high efficiency.
@@
In real-world scenarios, some noise may not be effectively handled by source separation, the ASR step may introduce errors, and some raw speech data may be of low quality [12]. Therefore, to ensure the quality of the resulting dataset, we apply the following filtering criteria. [^9] Please note that the filtering criteria can be adjusted to fit the specific needs of different use cases. Firstly, we utilize the language identification results from the Whisper model in the ASR step. We discard speech data that are not predicted to belong to our target languages (En, Zh, De, Fr, Ja, Ko) or have a language identification confidence lower than 80%. Secondly, we use the DNSMOS P.835 OVRL score [37] (hereafter referred to as DNSMOS score for brevity) to estimate the overall speech quality, preserving only those speech data with a score higher than 3.0. Finally, for each raw speech data, we compute the average character duration over its corresponding segments. We consider segments with an average phone duration outside 1.5 times the interquartile range (IQR) above the third quartile or below the first quartile as outliers and discard the speech data for these segments. After filtering, we obtain the resulting high-quality dataset suitable for training the speech generation model.
@@
TABLE II: Statistics of 600 hours of in-the-wild raw speech data processed by Emilia-Pipe.
Dataset Duration (s) DNSMOS Total Duration (hours)
min max avg ± std min max avg ± std
Raw 9.22 18,056.98 1,572.53 ± 1,966.66 1.08 3.51 2.50 ± 0.62 598.87 (100.00%)
Processed w/o Filtering 1.00 30.00 7.18 ± 5.06 0.91 3.67 2.86 ± 0.51 340.54 (56.86%)
Processed 3.00 30.00 8.98 ± 4.99 3.00 3.67 3.26 ± 0.14 176.22 (29.43%)
To validate the effectiveness of Emilia-Pipe, we randomly sample a subset of raw speech data, approximately 600 hours, and use Emilia-Pipe to process this subset to evaluate its effectiveness and efficiency. The evaluation is conducted on an independent server with eight NVIDIA RTX 4090 GPUs. The entire processing time takes about 3.99 hours. [^10] Please note that the processing speed may be subject to hardware environments and data characteristics. The reported figure is for reference. Table II shows the processing results of Emilia-Pipe on this subset. The raw data has a wide range of audio durations from 9.22 to 18,056.98 seconds, with an average of 1,572.53 seconds and high variability. The DNSMOS scores range from 1.08 to 3.51, with an average of 2.50, indicating varied overall quality. After filtering, the total duration for the resulting data is further reduced to 176.22 hours, retaining 29.43% of the raw speech data, and the average DNSMOS score significantly improves to 3.26 with minimal variability, indicating that Emilia-Pipe can effectively transform in-the-wild speech data into high-quality training data for speech generation. Additionally, for processing this subset, Emilia-Pipe processes about 2.50 hours of data every minute, demonstrating our Emilia-Pipe significantly exceeds real-time standards, making it ideal for preprocessing extensive speech data and scaling up the training dataset.

### v2

@@ Abstract
Recent advancements in speech generation have been driven by large-scale training datasets. However, current models struggle to capture the spontaneity and variability inherent in real-world human speech, as they are primarily trained on audio-book datasets limited to formal, read-aloud speaking styles. To address this limitation, we introduce Emilia-Pipe, an open-source preprocessing pipeline designed to extract high-quality training data from valuable yet under-explored in-the-wild sources that capture spontaneous human speech in real-world contexts. Using Emilia-Pipe, we construct Emilia, which comprises over 101k hours of speech across six languages: English, Chinese, German, French, Japanese, and Korean. Furthermore, we expand Emilia to Emilia-Large, a dataset exceeding 216k hours, making it one of the largest open-source speech generation resources available. Extensive experiments show that Emilia-trained models produce markedly more spontaneous, human-like speech than those trained on traditional audio-book datasets, while matching their intelligibility. These models better capture diverse speaker timbres and the full spectrum of real-world conversational styles. Our work also highlights the importance of scaling dataset size for advancing speech generation performance and validates the effectiveness of Emilia for both multilingual and crosslingual speech generation tasks.
@@
In-the-wild Speech Generation Dataset
@@ I Introduction
In recent years, (zero-shot) speech generation research has witnessed significant advancements [1, 2, 3, 4, 5, 6], with multiple models driven by large-scale training datasets. These advancements lead to improved voice quality, timbre similarity, and naturalness [7]. Nevertheless, the generated speech still falls short of replicating the spontaneity and variability characteristic of real-world human speech [2, 8].
@@
A primary factor for this limitation is the reliance of current speech generation models on datasets derived from audio-books [9, 10]. Such datasets predominantly feature formal, read-aloud speaking styles, which contrast with the diverse and spontaneous nature of human speech in casual or conversational settings. Such real-world speech is characterized by a wide range of phenomena, including breathing sounds, pauses, stuttering, repetitions, variations in speaking rate and emotions. Consequently, there remains a significant research gap for a dataset that encompasses a wider spectrum of speaking styles for speech generation model training.
@@
However, directly utilizing in-the-wild speech data presents significant challenges due to variations in quality, such as frequent background noise or music, reverberation, overlapping speakers within a single sample, inconsistent speech lengths, and the absence of essential annotations like text transcriptions [11, 12]. Training speech generation models on such unprocessed raw data can result in degraded performance [11, 13, 14]. While previous studies [11, 12] have proposed automatic preprocessing pipelines to mitigate these issues, they heavily depend on proprietary models, which significantly limits their accessibility for the broader research community. Moreover, these pipelines are typically restricted to monolingual (i.e., Chinese-only) speech data, making them unsuitable for processing the multilingual speech data available in the wild. Finally, the computational efficiency of these approaches also remains undocumented, raising concerns about their practicality for building large-scale speech generation datasets. Therefore, there is an urgent need for an effective and open-source preprocessing pipeline that can efficiently handle multilingual in-the-wild speech data and enable large-scale dataset construction.
@@
In response, we present Emilia-Pipe, the first open-source preprocessing pipeline specifically designed to leverage valuable yet underexplored in-the-wild multilingual speech data for constructing high-quality training datasets for spontaneous and human-like speech generation models. Emilia-Pipe comprises six core preprocessing steps: standardization, source separation, speaker diarization, fine-grained segmentation by voice activity detection (VAD), automated speech recognition (ASR), and filtering. In addition, Emilia-Pipe incorporates extensive engineering optimizations to enhance both robustness and efficiency. These features position Emilia-Pipe as an effective and efficient tool for building large-scale, multilingual speech datasets.
Utilizing Emilia-Pipe, we introduce Emilia, the first multilingual speech generation dataset constructed from in-the-wild speech data. Emilia comprises over 101k hours of 24 kHz speech across six languages: English (En), Chinese (Zh), German (De), French (Fr), Japanese (Ja), and Korean (Ko). To further scale up our dataset, we present Emilia-Large, which expands the total duration to 216k hours. Table I compares Emilia and Emilia-Large with several existing datasets.
The main advantages of our proposed Emilia and Emilia-Large datasets are summarized below.
@@
  • Extensive: Emilia contains over 101k hours of speech data, and the Emilia-Large variant expands the size to 216k hours, which is one of the largest open-source speech generation datasets.
@@
  • Multilingual: The Emilia and Emilia-Large datasets cover six languages, supporting the training of multilingual and crosslingual speech generation models.
@@
  • Diverse: Emilia and Emilia-Large distinguish themselves from prior datasets by centering on spontaneous, in-the-wild speech. This ensures coverage of a broad spectrum of speaking styles, which is essential for training next-generation speech generation models capable of producing natural, spontaneous, and human-like speech.
@@
  • Dynamic: The Emilia and Emilia-Large datasets uniquely feature an automatic and efficient processing pipeline, i.e., Emilia-Pipe, which enables seamless expansion in both dataset size and language coverage, significantly accelerating dataset construction.
@@
Libri-Heavy [21] Audio-book 50k En 16k
MLS [10] Audio-book 51k En/Fr/De/Nl/Es/It/Pt/Pl 16k
Libri-Light [9] Audio-book 60k En 16k
SeamlessAlign [22] In-the-wild 270k 76 Languages 16k
@@
  • Larger-scale Dataset: We expanded the initial Emilia dataset to create Emilia-Large, a dataset more than twice the size of its predecessor.
  • Comparative Analysis of Audio-book and In-the-Wild Datasets: We train identical models on traditional audio-book datasets and on the in-the-wild Emilia dataset, then compare their speech generation performance. Results demonstrate that models trained with in-the-wild data achieve significantly higher speaker similarity and naturalness in the synthesized speech.
  • Exploration of Data Scaling Laws in Speech Generation: We conducted experiments to investigate the effect of dataset size on speech generation performance. The results highlight the importance of scaling dataset sizes.
  • Multilingual and Crosslingual Effectiveness Analysis: We conduct experiments to validate the multilingual and crosslingual capabilities of Emilia, further demonstrating its applicability across six different languages.
Our code for Emilia-Pipe [^1] https://github.com/open-mmlab/Amphion/tree/main/preprocessors/Emilia as well as the Emilia and Emilia-Large datasets [^2] https://huggingface.co/datasets/amphion/Emilia-Dataset have been made publicly available to facilitate future research and ensure reproducibility. The remainder of this paper is organized as follows: Section II reviews related work. Section III describes the proposed Emilia-Pipe. Section IV details and analyzes our datasets. Experimental results are presented in Section V. Section VI concludes the paper.
@@ II-A Speech Generation Datasets
The size of datasets for speech generation has increased substantially over the years. Early datasets typically comprised tens of hours of speech. For example, LJSpeech [15] contains 24 hours of speech data from a single speaker. The VCTK dataset [16] includes 44 hours of speech data from 109 speakers, while AISHELL-3 [17] comprises approximately 85 hours of recordings from 218 speakers. Subsequently, larger-scale datasets have emerged to enable research in zero-shot speech generation. For instance, LibriTTS [18] contains 585.8 hours of speech data from audio-books, Later, the research community scaled up speech generation datasets to over 10k hours. Emerging datasets such as MLS [10] (51k hours), Libri-Light [9] (60k hours), and Libri-Heavy [21] (50k hours) significantly enhance the performance for zero-shot speech generation models. However, these datasets are derived from audio-books and thus primarily capture formal, read-aloud speaking style, limiting their effectiveness for training natural and spontaneous speech generation models [8, 24]. While large-scale in-the-wild corpora such as GigaSpeech [19], WenetSpeech [20], and SeamlessAlign [22] are available alternatives for broader speaking styles, their uncurated content and inconsistent speech quality can severely degrade generation performance if used without careful preprocessing [11, 13].
To address this, two previous works [12, 11] propose similar automated preprocessing pipelines for building speech generation datasets from in-the-wild data. However, as discussed earlier, these pipelines heavily rely on proprietary models and are restricted to Chinese-only speech data, with unknown computational efficiency. To bridge this gap, we design and open-source an efficient pipeline, Emilia-Pipe, which can rapidly process large-scale raw multilingual speech data to facilitate large-scale dataset construction.
@@ II-B Speech Generation Models
Traditional speech generation models, such as Tacotron [25], FastSpeech [26], are limited by smaller datasets with only tens of hours of speech data. Recent advancements in speech generation, driven by large-scale speech datasets, have significantly improved voice quality, timbre similarity, and naturalness in zero-shot speech generation using only a short reference speech sample of a few seconds. For example, VALLE [1], VoiceBox [3], and SoundStorm [4] use more than 50k hours of speech data for training. Subsequent models, such as NaturalSpeech3 [2], BaseTTS [5], and Seed-TTS [6], further expanded dataset sizes to over 100k hours. Notably, BaseTTS reported the “emergent abilities” of TTS models: as dataset sizes scaled, the models can render complex prosody patterns such as emotions based on textual cues without explicit labels. In this work, we introduce Emilia-Large, one of the largest open-source datasets for speech generation of 216k hours, to further advance the development of speech generation.
Figure 1: An overview of the Emilia-Pipe pipeline. It consists of six steps, namely, standardization, source separation, speaker diarization, fine-grained segmentation by voice activity detection (VAD), automated speech recognition (ASR), and filtering.
@@ III The Emilia-Pipe Processing Pipeline
As discussed above, source speech data in-the-wild need to be processed to be leveraged for training speech generation models. Therefore, we design an automatic preprocessing pipeline, Emilia-Pipe, for transforming in-the-wild multilingual speech data into high-quality training datasets. As illustrated in Fig. 1, Emilia-Pipe includes six steps: standardization, source separation, speaker diarization, fine-grained segmentation by voice activity detection (VAD), automated speech recognition (ASR), and filtering. This section details the six steps of Emilia-Pipe and evaluates its performance.
@@
The source speech data in-the-wild vary in encoding formats, sampling rates, and other characteristics. To standardize the collected data, we convert all samples to WAV files, set them to a mono channel, and resample to 24 kHz. We set the sample width to 16 bits and adjust the target decibels relative to full scale to -20 dBFS. The actual gain is constrained within -3 to 3 dB to ensure appropriate volume without distortion. Finally, we normalize the waveform by dividing each sample by the maximum amplitude, ensuring values range between -1 and 1. This step ensures a consistent data format for further processing.
@@
The source speech data in-the-wild often contain background music and noise, which negatively impact the performance of speech generation models [3, 13]. To address this issue, we employ source separation techniques to extract clean human vocals. Specifically, we utilize the open-source pre-trained Ultimate Vocal Remover model [27], UVR-MDX-Net Inst 3333https://github.com/TRvlvr/model_repo/releases/tag/all_public_uvr_models. This model achieves a high signal-to-distortion ratio (SDR) of 11.15 for vocal separation on the Synth MVSep dataset [28]. Using this model, we effectively separate human vocals for further processing.
@@
After extracting clean human vocals from the source speech data, we apply speaker diarization techniques to partition long-form speech data into multiple utterances based on the speaker. This process generates a series of speech segments, with each segment ideally containing only one speaker, ensuring compatibility with existing datasets for speech generation. To achieve this, we leverage the PyAnnote speaker diarization 3.1 pipeline. [^4] https://github.com/pyannote/pyannote-audio, which includes three core components: speaker segmentation, speaker embedding, and clustering, achieving state-of-the-art (SOTA) speaker diarization performance [29, 30]. The output is a list of temporal annotations indicating the start and end times of the single-speaker segments.
@@
Although the speaker diarization pipeline provides a coarse segmentation for the source speech data, the resulting utterances may still be too long for training speech generation models. To address this, we use a VAD model to further segment the utterances into smaller segments ranging from 3 to 30 seconds. Specifically, we concatenate consecutive chunks containing voice activity from the same speaker. We leverage the open-source library Silero-VAD [^5] https://github.com/snakers4/silero-vad, which achieves a ROC-AUC score of 0.99 on the LibriParty dataset. [^6] https://github.com/speechbrain/speechbrain/tree/develop/recipes/LibriParty/generate_dataset
@@
The absence of text transcriptions impedes the direct use of in-the-wild dataset for TTS. Therefore, we use ASR techniques to transcribe the segmented speech data. Considering the trade-off among speed, robustness, and accuracy, we employ the medium version of the Whisper model [31], a SOTA multilingual ASR model capable of robust speech translation and language identification. To further enhance efficiency, we leverage WhisperX [32], which builds on the faster-whisper backend [^7] https://github.com/SYSTRAN/faster-whisper and the CTranslate2 engine. [^8] https://github.com/OpenNMT/CTranslate2 This setup is up to four times faster than the official Whisper implementation while maintaining comparable accuracy and using less memory. Additionally, we omit the original model’s inherent VAD component by using the outputs in the last step to avoid redundant processing and develop batched inference to transcribe the speech data in parallel. These improvements allow our ASR step to achieve accurate text transcriptions for the speech data with high efficiency.
TABLE II: Processing results of Emilia-Pipe.
Processing Steps Duration (s) DNSMOS $\uparrow$ Clips Total Duration (hours)
min max avg ± std min max avg ± std
Source Speech 20.09 3596.27 547.80 ± 653.92 1.04 3.51 2.52 ± 0.68 4383 666.94 (100.00%)
+ Source Separation 20.09 3596.27 547.80 ± 653.92 0.96 3.54 2.88 ± 0.50 4383 666.94 (100.00%)
+ Speaker Diarization 0.02 1955.93 4.30 ± 10.73 0.64 3.69 2.50 ± 0.69 431876 515.50 (77.29%)
+ Fine-grained Segmentation by VAD 3.00 30.00 9.05 ± 5.25 0.80 3.69 2.92 ± 0.50 207709 522.03 (78.27%)
+ ASR 3.00 30.00 9.16 ± 5.25 0.97 3.68 2.98 ± 0.44 172923 439.80 (65.94%)
+ Filtering 3.00 30.00 9.68 ± 5.17 3.00 3.68 3.26 ± 0.14 96117 258.44 (38.75%)
Total Processing Time: 240.5 mins; Real-Time Factor (RTF): 0.006
@@
Real-world noises may not be completely handled by source separation, the ASR step may introduce errors, and some source speech data may be of low quality [11]. Therefore, to ensure the quality of the resulting dataset, we apply the following filtering criteria. [^9] Please note that the filtering criteria can be adjusted to fit the specific needs of different use cases. Firstly, we utilize the language identification results from the Whisper model in the ASR step. We discard speech data that are not predicted to belong to our target languages (En, Zh, De, Fr, Ja, Ko) or have a language identification confidence lower than 80%. Secondly, we use the DNSMOS P.835 OVRL score [33] (hereafter referred to as DNSMOS score for brevity) to estimate the overall speech quality. This non-intrusive metric reflects the overall quality of the speech data and is highly correlated with human ratings [33]. Following ITU-T P.835 [34], we retain only utterances whose DNSMOS $\geq$ 3.0, the lowest “fair” quality anchor, thereby guaranteeing that every training sample meets a minimum bar accepted for high-fidelity speech synthesis. Finally, for each source speech data, we compute the average character duration over its corresponding segments. We consider segments with an average phone duration outside 1.5 times the interquartile range (IQR) above the third quartile or below the first quartile as outliers and discard the speech data for these segments. After filtering, we obtain the resulting dataset.
@@
To evaluate the effectiveness of Emilia-Pipe, we processed 666.94 hours of randomly selected source speech data through the pipeline. Table II illustrates the impact of each processing step within Emilia-Pipe.
Before processing, the raw source speech data is highly variable, with clip durations ranging from 20.09 to 3596.27 seconds (averaging 547.80 seconds $\pm$ 653.92 seconds), and DNSMOS scores spanning 1.04 to 3.51 (averaging 2.52 $\pm$ 0.68), reflecting varied speech quality in-the-wild. Applying source separation preserves the total duration but improves the average DNSMOS score to 2.88 $\pm$ 0.50. Speaker diarization segments multi-speaker audio, increasing the number of clips to 431,876 while reducing the total duration to 515.50 hours (77.29% of the original). Subsequent VAD steps further refine the dataset, constraining all segment durations to between 3 and 30 seconds. The final filtering step yields 96,117 high-quality clips totaling 258.44 hours (38.75% of the original data), with a notably improved DNSMOS score of 3.26 $\pm$ 0.14, indicating low variability and superior speech quality. These results demonstrate Emilia-Pipe’s capability to convert raw speech data in the wild into a high-quality dataset for speech generation.
This experiment was conducted on a server equipped with eight NVIDIA RTX 4090 GPUs, running eight independent processes. [^10] Processing speed may vary depending on hardware and data characteristics; reported figures are provided for reference. On this server, processing the 666.94 hours of source speech data took 240.5 minutes, resulting in a Real-Time Factor (RTF) of approximately 0.006 (i.e., 1 second of source speech data processed in 0.006 seconds). This high efficiency underscores Emilia-Pipe’s suitability for large-scale speech dataset preprocessing, enabling fast training data preparation for speech generation models.